What is the approximate area of a segment of a circle with a radius 12 meters if the length of the chord is 20 meters?
Answer by RAY100(1637)

draw a rough sketch, circle of radius 12, with perpendicular chord of 20 bisected by
a radius. Let angle a = arc sin 10/12 = 56.4 deg
.
2a = 112. 885 deg,,,,which is central angle of segment
.
Area of circle = pi * r^2 = pi * 12 ^2 = 542.389
.
Area of segment is 112 / 360 * 542 = 142 meters squared